OpenStudy (tanabugg):

Pyramid base= 10 cm Height= 12 cm What is the slant height for the given pyramid? Answers: 7cm 11cm 13cm 16cm

OpenStudy (tanabugg):

@rebeccaxhawaii

OpenStudy (tanabugg):

@Danny_Boy420

rebeccaxhawaii (rebeccaxhawaii):

how do you calculate

OpenStudy (tanabugg):

I don't know...

rebeccaxhawaii (rebeccaxhawaii):

Find the slant height of the square pyramid. The slant height is the hypotenuse of the right triangle formed by the height and half the base length. Use the Pythagorean Theorem.

OpenStudy (anonymous):

im looking it up now

OpenStudy (tanabugg):

Oh so I just need the width and height

OpenStudy (welshfella):

is this a square pyramid?

OpenStudy (welshfella):

if so we can slice it down the middle and you;ll get the right angled trianggle 

OpenStudy (welshfella):

Find X

OpenStudy (welshfella):

the base is 5 because the length of the side of base is 10

OpenStudy (tanabugg):

12*12= 144 5*5= 25 144*25= 3600

OpenStudy (tanabugg):

the square root of 3600 is 60

OpenStudy (welshfella):

no its 144 + 25

OpenStudy (tanabugg):

Oh it's 13 then

OpenStudy (welshfella):

yes

OpenStudy (tanabugg):

Thanks!

OpenStudy (welshfella):

yw